OpenStudy (anonymous):

Find the ratio of the area of triangle XBY to the area of triangle ABC for the given measurements, if XY l l AC XY = 2, AC = 3 CLICK THE PIC BELOW

OpenStudy (anonymous):

where is the pic?

OpenStudy (anonymous):

Interesting question. The ratio is \[(AC/XY)^{2}\] = \[(3/2)^{2}\] = 9/4 Why? Area is (base x height)/2. Finding height of isosceles triangle... cut the triangle in two parts. You will find two pitagoric triangles. Formula for this case is h\[h^{2} = side^{2} - (base/2)^{2}\] base/2 because new triangles has this base. Solving, \[h= \sqrt{lado^{2}-(base/2)^{2}}\] Area is \[1/2*base*\sqrt{lado^{2}-(base/2)^{2}}\] When you increase sidesproportionaly the new area is \[1/2*3/2*base*\sqrt{(3/2lado)^{2}-(3/2*base/2)^{2}} =\] \[1/2*3/2*base*3/2*\sqrt{(lado)^{2}-(base/2)^{2}}\], jumping out 3/2 of square root. The diference betwenn cases is 3/2 and another 3/2. Finally 9/4.